The following information was drawn from the year-end balance sheets of Munoz River, Inc.:

Account Title 2017 2016
Bonds payable $ 785,000 $ 980,000
Common stock 201,000 126,000
Treasury stock 26,500 6,500
Retained earnings 86,400 59,900

Additional information regarding transactions occurring during 2017:

  1. Munoz River, Inc., issued $48,300 of bonds during 2017. The bonds were issued at face value. All bonds retired were retired at face value.

  2. Common stock did not have a par value.

  3. Munoz River, Inc., uses the cost method to account for treasury stock.

  4. The amount of net income shown on the 2017 income statement was $33,200.

Required

  1. Determine the amount of cash flow for the retirement of bonds that should appear on the 2017 statement of cash flows.

  2. Determine the amount of cash flow from the issue of common stock that should appear on the 2017 statement of cash flows.

  3. Determine the amount of cash flow for the purchase of treasury stock that should appear on the 2017 statement of cash flows.

  4. Determine the amount of cash flow for the payment of dividends that should appear on the 2017 statement of cash flows.

  5. Prepare the financing activities section of the 2017 statement of cash flows.
Answer #1

a) Cash flow for the retirement of bonds = 980000+48300-785000 = $243300

b) Issue common Stock = 201000-126000 = $75000

c) Cash flow from purchase of treasury stock = 26500-6500 = $20000

d) Cash flow from paying dividend = 59900+33200-86400 = 6700

e) Financing activities

Issue Bonds payable 48300
Retirement of bonds -243300
Issue common Stock 75000
Purchase treasury stock -20000
Dividend paid -6700
Net cash used in financing activities -146700